# High-Dimensional Gaussian Process Regression with Soft Kernel Interpolation

### v2

@@ 5 Experiments
We compare the performance of SoftKIs against popular scalable GPs on selected data sets from the UCI data set repository (Kelly et al., 2017), a common GP benchmark (Section 5.1). Next, we test SoftKIs on high-dimensional molecule data sets from the domain of computational chemistry (Section 5.2). Finally, we explore the interpolation points learned by SoftKI (Section 5.3).
@@ 5.1 Benchmark on UCI Regression
$n$ $d$ SoftKI $m{=}512$ SGPR $m{=}512$ SVGP $m{=}1024$
3droad 391386 3 0.583 $\pm$ 0.01 - 0.389 $\pm$ 0.001
Kin40k 36000 8 0.169 $\pm$ 0.008 0.177 $\pm$ 0.005 0.165 $\pm$ 0.005
Protein 41157 9 0.596 $\pm$ 0.016 0.601 $\pm$ 0.015 0.607 $\pm$ 0.012
Houseelectric 1844352 11 0.047 $\pm$ 0.001 - 0.047 $\pm$ 0.0
Bike 15641 17 0.062 $\pm$ 0.001 0.099 $\pm$ 0.003 0.084 $\pm$ 0.006
Elevators 14939 18 0.36 $\pm$ 0.006 0.393 $\pm$ 0.007 0.384 $\pm$ 0.008
Keggdirected 43944 20 0.08 $\pm$ 0.005 0.251 $\pm$ 0.13 0.082 $\pm$ 0.004
Pol 13500 26 0.075 $\pm$ 0.002 0.108 $\pm$ 0.001 0.122 $\pm$ 0.002
Keggundirected 57247 27 0.115 $\pm$ 0.004 0.132 $\pm$ 0.021 0.121 $\pm$ 0.007
Buzz 524925 77 0.24 $\pm$ 0.001 - 0.25 $\pm$ 0.002
Song 270000 90 0.777 $\pm$ 0.004 - 0.794 $\pm$ 0.006
Slice 48150 385 0.022 $\pm$ 0.006 0.435 $\pm$ 0.003 0.082 $\pm$ 0.001
Table 1: Test RMSE on UCI datasets. Best results are bolded.
$n$ $d$ SoftKI $m{=}512$ SGPR $m{=}512$ SVGP $m{=}1024$
3droad 391386 3 0.953 $\pm$ 0.041 - 0.597 $\pm$ 0.008
Kin40k 36000 8 0.055 $\pm$ 0.043 -0.105 $\pm$ 0.01 -0.082 $\pm$ 0.007
Protein 41157 9 0.905 $\pm$ 0.026 1.029 $\pm$ 0.01 1.047 $\pm$ 0.009
Houseelectric 1844352 11 1.274 $\pm$ 1.425 - -1.492 $\pm$ 0.007
Bike 15641 17 -0.379 $\pm$ 0.016 -0.322 $\pm$ 0.01 -0.68 $\pm$ 0.017
Elevators 14939 18 0.406 $\pm$ 0.021 0.581 $\pm$ 0.008 0.586 $\pm$ 0.009
Keggdirected 43944 20 0.421 $\pm$ 0.063 2.355 $\pm$ 3.875 -0.934 $\pm$ 0.017
Pol 13500 26 -0.71 $\pm$ 0.028 -0.538 $\pm$ 0.003 -0.394 $\pm$ 0.012
Keggundirected 57247 27 0.302 $\pm$ 0.136 -0.56 $\pm$ 0.108 -0.59 $\pm$ 0.021
Buzz 524925 77 0.276 $\pm$ 0.348 - 0.13 $\pm$ 0.008
Song 270000 90 1.179 $\pm$ 0.008 - 1.288 $\pm$ 0.002
Slice 48150 385 1.258 $\pm$ 0.149 1.295 $\pm$ 0.003 -0.662 $\pm$ 0.002
Table 2: Test NLL on UCI datasets. Best results are bolded.
Experiment details. For this experiment, we split the data set into $0.9$ for training and $0.1$ for testing. We standardize the data to have mean 00 and standard deviation $1$ using the training data set. We use a Matérn 3/2 kernel and a learnable output scale. We choose $m=512$ inducing points for SoftKI. Following Wang et al. (2019), we use $m=512$ for SGPR and $m=1024$ for SVGP. We learn model hyperparameters for SoftKI by maximizing Hutchinson’s pseudoloss, and for SGPR and SVGP by maximizing the ELBO.
@@ 5.2 High-Dimensional Molecule Dataset
Since SoftKI’s are performant on large dimensional data sets from the UCI repository, we also test the performance on molecular potential energy surface data on the MD22 (Chmiela et al., 2023) dataset. These are high-dimensional datasets that give the geometric coordinates of atomic nuclei in biomolecules and their respective energies.
$n$ $d$ SoftKI $m{=}512$ SGPR $m{=}512$ SVGP $m{=}1024$
Ac-ala3-nhme 76598 126 0.79 $\pm$ 0.01 0.847 $\pm$ 0.006 0.836 $\pm$ 0.005
Dha 62777 168 0.886 $\pm$ 0.012 0.897 $\pm$ 0.012 0.882 $\pm$ 0.01
Stachyose 24544 261 0.363 $\pm$ 0.012 0.638 $\pm$ 0.003 0.555 $\pm$ 0.001
At-at 18000 354 0.528 $\pm$ 0.011 0.589 $\pm$ 0.008 0.558 $\pm$ 0.009
At-at-cg-cg 9137 354 0.502 $\pm$ 0.007 0.563 $\pm$ 0.024 0.461 $\pm$ 0.026
Buckyball-catcher 5491 444 0.153 $\pm$ 0.006 0.394 $\pm$ 0.011 0.307 $\pm$ 0.015
Double-walled-nanotube 4528 1110 0.031 $\pm$ 0.001 1.001 $\pm$ 0.048 0.045 $\pm$ 0.0
Table 3: Test RMSE on MD22 datasets. Best results are bolded.
$n$ $d$ SoftKI $m{=}512$ SGPR $m{=}512$ SVGP $m{=}1024$
Ac-ala3-nhme 76598 126 1.188 $\pm$ 0.014 1.364 $\pm$ 0.003 1.356 $\pm$ 0.002
Dha 62777 168 1.299 $\pm$ 0.014 1.418 $\pm$ 0.006 1.409 $\pm$ 0.007
Stachyose 24544 261 0.522 $\pm$ 0.059 1.118 $\pm$ 0.003 1.012 $\pm$ 0.003
At-at 18000 354 0.795 $\pm$ 0.007 1.037 $\pm$ 0.004 1.003 $\pm$ 0.008
At-at-cg-cg 9137 354 0.742 $\pm$ 0.01 0.987 $\pm$ 0.017 0.8 $\pm$ 0.016
Buckyball-catcher 5491 444 -0.278 $\pm$ 0.061 0.69 $\pm$ 0.009 0.449 $\pm$ 0.015
Double-walled-nanotube 4528 1110 -1.551 $\pm$ 0.049 1.52 $\pm$ 0.021 -0.954 $\pm$ 0.009
Table 4: Test NLL on MD22 datasets. Best results are bolded.
@@
Experiment details. For consistency, we keep the experimental setup the same as for UCI regression but up the number of epochs of training to 200 due to slower convergence on these datasets. We still use the Matérn kernel and not a molecule-specific kernel that incorporates additional information such as the atomic number of each atom (e.g., a Hydrogen atom) or invariances (e.g., rotational invariance). We standardize the data set to have mean 00 and standard deviation $1$. We note that in an actual application of GP regression to this setting, we may only opt to center the targets to be mean 00. This is because energy is a relative number that can be arbitrarily shifted, whereas scaling the distances between atoms will affect the physics.
@@
Results. Table 4 and Table 4 compares the test RMSE and test NLL of various GP models trained on MD22. We see that SoftKI is a competitive method on these datasets, especially on the test NLL. GPs that fit forces have been successfully applied to fit such data sets to chemical accuracy. In our case, we do not fit derivative information. It would be an interesting direction of future work to extend SoftKI to handle derivatives.
5.3 Learned Interpolation Points
Figure 3: (Interpolation vs. Inducing points–High Dimensional Setting): In this comparison the slice ($d=385$) dataset’s PCA embedding is presented, along with a kernel density estimate of the embedded training data to more clearly highlight the placement of the learned interpolation points learned by SoftKI, and inducing points learned by SGPR and SVGP relative to the distribution of the data. Importantly note that in this example SVGP is run with twice as many ($1024$) inducing points compared to the other methods. Appendix C.4 contains the PCA analysis for the remaining UCI datasets.
Because a SoftKI’s predictive performance is determined by where it places its learned interpolation points, we examine their arrangement for the slice dataset in Figure 3. We apply a two-component PCA embedding to both the training dataset and the interpolation/inducing points $\mathbf{z}$ learned by SoftKI, SGPR, and SVGP. To assess how effectively each method captures the overall data distribution, we plot the PCA embeddings of the inducing points overlaid on the kernel density estimation (KDE) contour levels of the training points in the PCA space, along with the marginal distributions along each axis.
In general, we see that the interpolation point locations learned by SoftKI are more spread out compared to the inducing points learned by SGPR and SVGP. We believe that this is because the prior on inducing variables linking the locations of inducing points to the data in SGPR and SVGP are normally distributed whereas there are no such assumptions in the case of SoftKI. Moreover, we optimize a stabilized MLL that is related to an exact GP’s MLL as opposed to an ELBO. Thus, while there are similarities in that both SoftKI and SGPR begin with a Nyström approximation of the covariance matrix, interpolation points and inducing points are subtly different. Additional interpolation point analysis, including PCA analysis for other datasets in the UCI repository, are included in Appendix C.4.
@@ 6 Conclusion
In this paper, we introduce SoftKI, an approximate GP designed for regression on large and high-dimensional datasets. SoftKI combines aspects of SKI and inducing points methods to retain the benefits of kernel interpolation while also scaling to higher dimensional datasets. We have tested SoftKI on a variety of datasets and shown that it is possible to perform kernel interpolation in high dimensional spaces in a way that is competitive with other approximate GP abstractions that leverage inducing points. Moreover, we find that the interpolation points learned by SoftKI are structurally different from the inducing points learned in SGRP and SVGP. This suggests that methods that learn interpolation points provide another promising path for constructing approximate GPs. Additionally, exploring methods to enforce stricter sparsity in the interpolation matrices, such as through thresholding, could enable the use of sparse matrix data structures further improving the cost of inference.
Acknowledgements Chris Camaño acknowledges support by the National Science Foundation Graduate Research Fellowship under Grant No. 2139433. Additionally, we would like to extend our appreciation to Ethan N. Epperly for his thoughtful comments and advice on this work.
@@ Appendix A SoftKI Posterior Predictive Derivation
$p(f_{\mathbf{z}}(*)\,|\,y)=\mathcal{N}(\mathbf{\widehat{K}_{*\mathbf{z}}}\mathbf{\widehat{C}}^{-1}\mathbf{\widehat{K}_{\mathbf{z}\mathbf{x}}}\mathbf{\Lambda}^{-1}\mathbf{y},\mathbf{K}_{**}^{\text{SoftKI}}-\mathbf{K}_{*\mathbf{x}}^{\text{SoftKI}}(\mathbf{\Lambda}^{-1}-\mathbf{\Lambda}^{-1}\mathbf{\widehat{K}_{\mathbf{x}\mathbf{z}}}\mathbf{\widehat{C}}^{-1}\mathbf{\widehat{K}_{\mathbf{z}\mathbf{x}}}\mathbf{\Lambda}^{-1})\mathbf{K}_{\mathbf{x}*}^{\text{SoftKI}})$ (23)
@@
can be solved for $\alpha_{*}$ in the same way we solved for the posterior mean (with $\mathbf{y}$ instead of $\mathbf{K}_{\mathbf{x}*}^{\text{SoftKI}}$). Since this depends on the inference point *, we cannot precompute the result ahead of time as we could with the posterior mean. Nevertheless, the intermediate results of the QR decomposition can be computed once during posterior mean inference and reused for the covariance prediction. Thus, the time complexity of posterior covariance inference is $\mathcal{O}(m^{2}n)$ per a test point.
Appendix B Algorithmic Choices
In this section, we examine various design choices we made for SoftKI. First, we discuss the rationale behind using a stochastic trace estimation scheme for computing the MLL during stochastic optimization (Section B.1). Next, we motivate choice of linear solver, comparing the use alternative linear solvers for the posterior inference procedure described in Section 4.3.
B.1 Effect of Hutchinson’s Pseudoloss
Figure 4: Comparison of Ricker wavelet surface reconstruction and interpolation points learned using the SoftKI MLL and Hutchinson’s pseudoloss with varying numbers of probes.
UCI Dataset Test RMSE Test NLL
dataset n d $\log\hat{p}(\mathbf{y}\,|\,\mathbf{x};\mathbf{\theta})$ $\log\tilde{p}(\mathbf{y}\,|\,\mathbf{x};\mathbf{\theta})$ $\log\hat{p}(\mathbf{y}\,|\,\mathbf{x};\mathbf{\theta})$ $\log\tilde{p}(\mathbf{y}\,|\,\mathbf{x};\mathbf{\theta})$
3droad 391386 3 0.583 $\pm$ 0.01 0.541 $\pm$ 0.005 0.953 $\pm$ 0.041 1.004 $\pm$ 0.175
Kin40k 36000 8 0.169 $\pm$ 0.008 0.183 $\pm$ 0.006 0.055 $\pm$ 0.043 0.074 $\pm$ 0.026
Protein 41157 9 0.596 $\pm$ 0.016 0.602 $\pm$ 0.016 0.905 $\pm$ 0.026 0.914 $\pm$ 0.029
Houseelectric 1844352 11 0.047 $\pm$ 0.0 0.049 $\pm$ 0.002 0.451 $\pm$ 0.029 1.275 $\pm$ 1.103
Bike 15641 17 0.062 $\pm$ 0.001 0.073 $\pm$ 0.005 -0.379 $\pm$ 0.016 -0.07 $\pm$ 0.216
Elevators 14939 18 0.36 $\pm$ 0.006 0.361 $\pm$ 0.007 0.406 $\pm$ 0.021 0.405 $\pm$ 0.024
Keggdirected 43944 20 0.08 $\pm$ 0.005 0.082 $\pm$ 0.004 0.421 $\pm$ 0.063 0.5 $\pm$ 0.232
Pol 13500 26 0.075 $\pm$ 0.002 0.084 $\pm$ 0.001 -0.71 $\pm$ 0.028 -0.744 $\pm$ 0.07
Keggundirected 57247 27 0.115 $\pm$ 0.004 0.118 $\pm$ 0.004 0.302 $\pm$ 0.136 0.315 $\pm$ 0.017
Buzz 524925 77 0.24 $\pm$ 0.001 0.24 $\pm$ 0.0 0.276 $\pm$ 0.348 0.29 $\pm$ 0.33
Song 270000 90 0.777 $\pm$ 0.004 0.777 $\pm$ 0.004 1.179 $\pm$ 0.008 1.178 $\pm$ 0.007
Slice 48150 385 0.022 $\pm$ 0.006 0.045 $\pm$ 0.004 1.258 $\pm$ 0.149 0.502 $\pm$ 0.137
Table 5: Effect on SoftKI RMSE and NLL using $\log\hat{p}(\mathbf{y}\,|\,\mathbf{x};\mathbf{\theta})$ vs. $\log\tilde{p}(\mathbf{y}\,|\,\mathbf{x};\mathbf{\theta})$.
In Section 4.2, we advocated for the use of Hutchinson’s pseudoloss to overcome numerical stability issues that arise when calculating the exact MLL. This adjustment is specifically to address situations where the matrix $\mathbf{K}_{\mathbf{z}\mathbf{z}}$ is not positive semi-definite (PSD) in float32 precision. In other approximate GP model using gpytorch, this challenge is typically met by performing a Cholesky decomposition followed by an efficient low-rank computation of the log determinant through the matrix determinant lemma. In our experience even when using versions of the Cholesky decomposition that add additional jitter along the diagonal there are still situations where $\mathbf{K}_{\mathbf{z}\mathbf{z}}$ can be poorly conditioned, particularity when $n$ is large. We conjecture that this behavior originates from situations arising where the learned interpolation points of a SoftKI coincide at similar positions driving the effective rank of $\mathbf{K}_{\mathbf{z}\mathbf{z}}$ down. In these situations, Hutchinson’s pseudoloss offers a more stable alternative because it does not directly rely on the matrix being invertible.
Figure 4 compares the results obtained with the SoftKI MLL and Hutchinson’s pseudoloss for varying numbers of probes on the Ricker wavelet surface. In Table 5 we replicate the UCI experiment of Section 5.1 using our stabilized MLL $\log\hat{p}(\mathbf{y}\,|\,\mathbf{x};\mathbf{\theta})$ and Hutchinson’s pseudoloss $\log\tilde{p}(\mathbf{y}\,|\,\mathbf{x};\mathbf{\theta})$. For most situations, the stabilized MLL and Hutchinson’s pseudoloss give similar results, indicating that the MLL is largely used. The results differ on some datasets that produce noisy $\mathbf{K}_{**}^{\text{SoftKI}}+\mathbf{\Lambda}$ which can cause the Cholesky decomposition to fail.
B.2 Alternative Methods for Posterior Inference

### v3

@@ 4 Experiments
We compare the performance of SoftKIs against popular scalable GPs on selected data sets from the UCI data set repository (Kelly et al., 2017), a common GP benchmark (Section 4.1). Next, we test SoftKIs on high-dimensional molecule data sets from the domain of computational chemistry (Section 4.2). Finally, we explore the numerical stability of SoftKI (Section 4.3).Our implementation can be found at: https://github.com/danehuang/softki_gp_kit.
Dataset $n$ $d$ Exact GP SoftKI $m=512$ SGPR $m=512$ SVGP $m=1024$
3droad 391386 3 -- $0.583\pm 0.010$ -- $\mathbf{0.389}\pm 0.001$
Kin40k 36000 8 $0.039\pm 0.001$ $0.169\pm 0.008$ $0.177\pm 0.004$ $\mathbf{0.165}\pm 0.005$
Protein 41157 9 $0.044\pm 0.000$ $\mathbf{0.596}\pm 0.016$ $0.602\pm 0.015$ $0.607\pm 0.012$
Houseelectric 1844352 11 -- $\mathbf{0.047\pm 0.001}$ -- $\mathbf{0.047}\pm 0.000$
[…]
Elevators 14939 18 $0.108\pm 0.014$ $\mathbf{0.360}\pm 0.006$ $0.395\pm 0.005$ $0.384\pm 0.008$
[…]
Keggundirected 57247 27 $0.110\pm 0.004$ $\mathbf{0.115}\pm 0.004$ $0.128\pm 0.009$ $0.121\pm 0.007$
Buzz 524925 77 -- $\mathbf{0.240}\pm 0.001$ -- $0.250\pm 0.002$
Song 270000 90 -- $\mathbf{0.777}\pm 0.004$ -- $0.794\pm 0.006$
Slice 48150 385 -- $\mathbf{0.022}\pm 0.006$ $520\pm 0.001 0.0$$082\pm 0.001$
Table 2: Test RMSE on UCI datasets. Best results obtained by an approximate GP are bolded. Exact GP included for reference. Entries marked "--" indicate datasets that triggered out of memory errors.
Dataset $n$ $d$ Exact GP SoftKI $m=512$ SGPR $m=512$ SVGP $m=1024$
3droad 391386 3 -- $0.953\pm 0.041$ -- $\mathbf{0.597}\pm 0.008$
Kin40k 36000 8 $1.636\pm 0.006$ $0.055\pm 0.043$ $\mathbf{-0.104}\pm 0.007$ $0.082\pm 0.007$
Protein 41157 9 $0.904\pm 0.008$ $\mathbf{0.905}\pm 0.026$ $1.031\pm 0.009$ $1.047\pm 0.009$
Houseelectric 1844352 11 -- $1.274\pm 1.425$ -- $\mathbf{-1.492}\pm 0.007$
Bike 15641 17 $1.662\pm 0.065$ $0.379\pm 0.016$ $0.296\pm 0.011$ $\mathbf{-0.680}\pm 0.017$
Elevators 14939 18 $0.590\pm 0.075$ $\mathbf{0.406}\pm 0.021$ $0.587\pm 0.006$ $0.586\pm 0.009$
Keggdirected 43944 20 nan $0.421\pm 0.063$ $0.896\pm 0.041$ $\mathbf{-0.934}\pm 0.017$
Pol 13500 26 $1.750\pm 0.020$ $\mathbf{-0.710}\pm 0.028$ $0.428\pm 0.002$ $0.394\pm 0.012$
Keggundirected 57247 27 nan $0.302\pm 0.136$ $\mathbf{-0.595}\pm 0.039$ $0.590\pm 0.021$
Buzz 524925 77 -- $0.276\pm 0.348$ -- $\mathbf{0.130}\pm 0.008$
Song 270000 90 -- $\mathbf{1.179}\pm 0.008$ -- $1.288\pm 0.002$
Slice 48150 385 -- $1.258\pm 0.149$ $1.330\pm 0.002$ $\mathbf{-0.662}\pm 0.002$
Table 3: Test NLL on UCI datasets. Best approximate results obtained by an approximate GP are bolded. Exact GP included for reference. Entries marked "--" indicate datasets that triggered out of memory errors. Entries marked nan indicate that the computations were unstable.
@@ 4.1 Benchmark on UCI Regression
Experiment details. For this experiment, we split the data set into $0.9$ for training and $0.1$ for testing. We standardize the data to have mean 0 and standard deviation $1$ using the training data set. We use a Matérn 3/2 kernel and a learnable output scale for each dimension (ARD). We choose $m=512$ inducing points for SoftKI. Following Wang et al. (2019), we use $m=512$ for SGPR and $m=1024$ for SVGP. We learn model hyperparameters for SoftKI by maximizing Hutchinson’s pseudoloss, and for SGPR and SVGP by maximizing the ELBO. For reference, we also include results for a GP fit with block CG, termed an Exact GP. For Exact GP we use the KeOps (Charlier et al., 2021) library integration with GPyTorch to handle the additional scaling challenges of working with the full $n\times n$ kernel.
@@ 4.2 High-Dimensional Molecule Dataset
Since SoftKI’s are performant on high dimensional data sets from the UCI repository, we also test the performance on molecular potential energy surface data on the MD22 (Chmiela et al., 2023) dataset. These are high-dimensional datasets that give the geometric coordinates of atomic nuclei in biomolecules and their respective energies.
Dataset $n$ $d$ Exact GP SoftKI m=512m{=}512 SGPR m=512m{=}512 SVGP m=1024m{=}1024
Ac-ala3-nhme 76598 126 $0.017\pm 0.000$ $\mathbf{0.790}\pm 0.010$ $0.852\pm 0.005$ $0.836\pm 0.005$
Dha 62777 168 $0.016\pm 0.000$ $0.886\pm 0.012$ $0.897\pm 0.012$ $\mathbf{0.882}\pm 0.010$
Stachyose 24544 261 $0.020\pm 0.000$ $\mathbf{0.363}\pm 0.012$ $0.643\pm 0.002$ $0.555\pm 0.001$
At-at 18000 354 $0.020\pm 0.001$ $\mathbf{0.528}\pm 0.011$ $0.591\pm 0.011$ $0.558\pm 0.009$
At-at-cg-cg 9137 354 $0.021\pm 0.000$ $0.502\pm 0.007$ $0.565\pm 0.029$ $\mathbf{0.461}\pm 0.026$
Buckyball-catcher 5491 444 $0.018\pm 0.001$ $\mathbf{0.153}\pm 0.006$ $0.393\pm 0.008$ $0.307\pm 0.015$
[…]
Table 4: Test RMSE on MD22 datasets. Best results obtained by an approximate GP are bolded. Exact GP results are included for reference. Exact GPs are roughly $20-875\times$ slower than SoftKI (see Table 10), and thus, only run for 50 epochs.
Dataset $n$ $d$ Exact GP SoftKI m=512m{=}512 SGPR m=512m{=}512 SVGP m=1024m{=}1024
Ac-ala3-nhme 76598 126 $0.000\pm 0.000$ $\mathbf{1.188}\pm 0.014$ $1.367\pm 0.003$ $1.356\pm 0.002$
Dha 62777 168 $0.975\pm 0.844$ $\mathbf{1.299}\pm 0.014$ $1.418\pm 0.006$ $1.409\pm 0.007$
Stachyose 24544 261 $1.330\pm 0.003$ $\mathbf{0.522}\pm 0.059$ $1.123\pm 0.002$ $1.012\pm 0.003$
At-at 18000 354 $1.418\pm 0.002$ $\mathbf{0.795}\pm 0.007$ $1.041\pm 0.006$ $1.003\pm 0.008$
At-at-cg-cg 9137 354 $1.426\pm 0.000$ $\mathbf{0.742}\pm 0.010$ $0.990\pm 0.024$ $0.800\pm 0.016$
Buckyball-catcher 5491 444 $1.567\pm 0.003$ $\mathbf{-0.278}\pm 0.061$ $0.689\pm 0.005$ $0.449\pm 0.015$
DW-nanotube 4528 1110 $1.992\pm 0.000$ $\mathbf{-1.551}\pm 0.049$ $1.520\pm 0.021$ $0.954\pm 0.009$
Table 5: Test NLL on MD22 datasets. Best results obtained by an approximate GP are bolded. Exact GP results are included for reference. Exact GPs are roughly $20-875\times$ slower than SoftKI (see Table 10), and thus, only run for 50 epochs.
@@
Experiment details. For consistency, we keep the experimental setup the same as for UCI regression but up the number of epochs of training to 200 due to slower convergence on these datasets. We still use the Matérn kernel and not a molecule-specific kernel that incorporates additional information such as the atomic number of each atom (e.g., a Hydrogen atom) or invariances (e.g., rotational invariance). We standardize the data set to have mean 0 and standard deviation 11. We note that in an actual application of GP regression to this setting, we may only opt to center the targets to be mean 0. This is because energy is a relative number that can be arbitrarily shifted, whereas scaling the distances between atoms will affect the physics. As before, we include results for Exact GPs as a reference. Since they are roughly $20$ to $875$ times slower than SoftKI, so we only run them for 50 epochs (see Table 10).
@@
Results. Table 5 and Table 5 compares the test RMSE and test NLL of various GP models trained on MD22. We see that SoftKI is a competitive approximate method on these datasets, especially on the test NLL. GPs that fit forces have been successfully applied to fit such data sets to chemical accuracy. In our case, we do not fit derivative information.
4.3 Numerical Stability
Test RMSE Test NLL
Dataset MLL Hutchinson Pseduoloss MLL Hutchinson Pseduoloss
3droad nan 0.541 $\pm$ 0.005 nan 1.004 $\pm$ 0.175
Kin40k 0.169 $\pm$ 0.008 0.183 $\pm$ 0.006 0.055 $\pm$ 0.043 0.07 $\pm$ 0.029
Protein 0.596 $\pm$ 0.016 0.602 $\pm$ 0.016 0.904 $\pm$ 0.03 0.914 $\pm$ 0.028
Houseelectric nan 0.049 $\pm$ 0.002 nan 1.275 $\pm$ 1.103
Bike nan 0.073 $\pm$ 0.005 nan -0.101 $\pm$ 0.179
Elevators 0.36 $\pm$ 0.007 0.36 $\pm$ 0.007 0.404 $\pm$ 0.024 0.404 $\pm$ 0.025
Keggdirected 0.08 $\pm$ 0.005 0.082 $\pm$ 0.004 0.293 $\pm$ 0.169 0.628 $\pm$ 0.289
Pol 0.075 $\pm$ 0.002 0.084 $\pm$ 0.001 -0.694 $\pm$ 0.06 -0.783 $\pm$ 0.033
Keggundirected 0.114 $\pm$ 0.005 0.118 $\pm$ 0.004 0.308 $\pm$ 0.025 0.291 $\pm$ 0.089
Buzz nan 0.24 $\pm$ 0.0 nan 0.599 $\pm$ 0.555
Song 0.777 $\pm$ 0.004 0.777 $\pm$ 0.004 1.179 $\pm$ 0.008 1.178 $\pm$ 0.007
Slice 0.022 $\pm$ 0.006 0.044 $\pm$ 0.005 1.328 $\pm$ 0.045 0.607 $\pm$ 0.058
Table 6: Comparison of SoftKI test RMSE and NLL when training with the SoftKI MLL $\log p(\mathbf{y}\mid\mathbf{x};\bm{\theta})$ versus the Hutchinson pseudoloss $\log\tilde{p}(\mathbf{y}\mid\mathbf{x};\bm{\theta})$. Best results are bolded. Entries labeled nan, represent instances where using the exact MLL resulted in numerical instability in SoftKI (see Section  4.3).
In Section 3.2, we advocated for the use of Hutchinson’s pseudoloss to overcome numerical stability issues that arise when calculating the SoftKI MLL. This adjustment is specifically to address situations where the matrix $\mathbf{K}_{\mathbf{x}_{b}\mathbf{x}_{b}}^{\text{SoftKI}}+\mathbf{\Lambda}$ is not positive semi-definite in float32 precision. In other approximate GP models implemented in GPyTorch, this challenge is typically addressed by performing a Cholesky decomposition followed by an efficient low-rank computation of the log determinant through the matrix determinant lemma.
In our experience, even when using versions of the Cholesky decomposition that add additional jitter along the diagonal there are still situations where $\mathbf{K}_{\mathbf{z}\mathbf{z}}$ can be poorly conditioned, particularly when $n$ is large. We conjecture that this behavior originates from situations where the learned interpolation points of a SoftKI coincide at similar positions, driving the effective rank of $\mathbf{K}_{\mathbf{z}\mathbf{z}}$ down. In these situations, Hutchinson’s pseudoloss offers a more stable alternative because it does not directly rely on the matrix being invertible.
In Table 6, we replicate the UCI experiment of Section 4.1 using the SoftKI MLL $\log p(\mathbf{y}\,|\,\mathbf{x};\mathbf{\theta})$ and Hutchinson’s pseudoloss $\log\tilde{p}(\mathbf{y}\,|\,\mathbf{x};\mathbf{\theta})$. For most datasets, optimizing with the SoftKI MLL produces better test RMSE. However, more numericaly instability is also encountered, with SoftKI failing on four datasets that we tested. In these cases, we find that Cholesky decomposition fails.
@@ 5 Conclusion
In this paper, we introduce SoftKI, an approximate GP designed for regression on large and high-dimensional datasets. SoftKI combines aspects of SKI and inducing points methods to retain the benefits of kernel interpolation while also scaling to higher dimensional datasets. We have tested SoftKI on a variety of datasets and shown that it is possible to perform kernel interpolation in high dimensional spaces in a way that is competitive with other approximate GP abstractions that leverage inducing points. Exploring methods that enforce stricter sparsity in the interpolation matrices, such as through thresholding, could enable the use of sparse matrix algebra further improving the cost of inference.
Acknowledgements Chris Camaño acknowledges support by the National Science Foundation Graduate Research Fellowship under Grant No. 2139433, and the Kortschak scholars program. Additionally, we would like to extend our appreciation to Ethan N. Epperly for his thoughtful comments and advice on this work.
@@ Appendix A SoftKI Posterior Predictive Derivation
$p(f(*)\,|\,\mathbf{x},\mathbf{y})=\mathcal{N}(\mathbf{\widehat{K}_{*\mathbf{z}}}\mathbf{\widehat{C}}^{-1}\mathbf{\widehat{K}_{\mathbf{z}\mathbf{x}}}\mathbf{\Lambda}^{-1}\mathbf{y},\mathbf{K}_{**}^{\text{SoftKI}}-\mathbf{K}_{*\mathbf{x}}^{\text{SoftKI}}(\mathbf{\Lambda}^{-1}-\mathbf{\Lambda}^{-1}\mathbf{\widehat{K}_{\mathbf{x}\mathbf{z}}}\mathbf{\widehat{C}}^{-1}\mathbf{\widehat{K}_{\mathbf{z}\mathbf{x}}}\mathbf{\Lambda}^{-1})\mathbf{K}_{\mathbf{x}*}^{\text{SoftKI}})$ (11)
@@
can be solved for $\alpha_{*}$ in the same way we solved for the posterior mean (with $\mathbf{y}$ instead of $\mathbf{K}_{\mathbf{x}*}^{\text{SoftKI}}$). Since this depends on the inference point *, we cannot precompute the result ahead of time as we could with the posterior mean. Nevertheless, the intermediate results of the QR decomposition can be computed once during posterior mean inference and reused for the covariance prediction. Thus, the time complexity of posterior covariance inference is $\mathcal{O}(nm^{2})$ per a test point.
Appendix B Alternative Methods for Posterior Inference